# Weak continuous measurements require more work than strong ones

### v2

@@ D.3 Work cost scaling
$W^{\text{strong}}=K\left[\frac{1}{2}\log{\left(2\pi\sigma^{2}\right)}+\frac{1}{2}\right],$ (73)
@@
$W^{\text{strong}}\propto K\left[l\log{\left(2\pi\epsilon\right)}+\frac{1}{2}\right].$ (74)
@@ Appendix E Hierarchy of information measures
For a mixture of states $\hat{\rho}_{\text{mixed}}=\sum_{x}p_{x}\hat{\rho}_{x}$, associated with random variable $X$ of probability distribution $p_{x}$, the Holevo bound states that the Holevo information $\chi(X)=S[\hat{\rho}_{\text{mixed}}]-\sum_{x}p_{x}S[\hat{\rho}_{x}]$ verifies [36]:
@@
We also note that the concavity of von Neumann entropy [36], together with the equalities $\hat{\rho}_{S}(t_{f})=\sum_{r}p_{r}\hat{\rho}_{S|r}$ and $\hat{\rho}_{S|r}=\frac{1}{p_{r}}\sum_{n}p_{n}p(r|n)\hat{\rho}_{S|n}$, imply:

### v3

@@ D.3 Work cost scaling
[…]
@@
[…]
@@ Appendix E Hierarchy of information measures
For a mixture of states $\hat{\rho}_{\text{mixed}}=\sum_{x}p_{x}\hat{\rho}_{x}$, associated with random variable $X$ of probability distribution $p_{x}$, the Holevo bound states that the Holevo information $\chi(X)=S[\hat{\rho}_{\text{mixed}}]-\sum_{x}p_{x}S[\hat{\rho}_{x}]$ verifies [Preskill16]:
@@
We also note that the concavity of von Neumann entropy [Preskill16], together with the equalities $\hat{\rho}_{S}(t_{f})=\sum_{r}p_{r}\hat{\rho}_{S|r}$ and $\hat{\rho}_{S|r}=\frac{1}{p_{r}}\sum_{n}p_{n}p(r|n)\hat{\rho}_{S|n}$, imply:
@@
Appendix F Pure dephasing, energy exchange, and the conditions for SBS formation
Consider an ancilla, modeled as a harmonic oscillator as in Sec. IV, coupled to several independent bosonic baths, indexed by $r$ and illustrated in Fig. 2. The total interaction Hamiltonian reads
$\hat{H}_{I}=\sum_{r}\hat{H}_{I,r},\qquad\hat{H}_{I,r}=\gamma_{r}\hat{a}^{\dagger}\hat{a}\otimes\hat{B}_{r},$ (83)
where $\gamma_{r}$ denotes the coupling strength and
$\hat{B}_{r}=\sum_{k}\hat{b}^{\dagger}_{k,r}\hat{b}_{k,r}.$ (84)
Each bath is assumed to start in a thermal state, which can be expressed in the coherent-state representation as
$\hat{\rho}_{B,r}=\int d^{2}\alpha_{r,l}P_{r,l}(\alpha_{r,l}){\left|{\alpha_{r,l}}\right\rangle}{\left\langle{\alpha_{r,l}}\right|}$ (85)
with the Glauber–Sudarshan distribution
$P(\alpha_{r,l})=\frac{1}{\pi\bar{n}}e^{-|\alpha_{r,l}|^{2}/\bar{n}},$ (86)
and mean occupation number
$\bar{n}=\frac{1}{e^{\beta\hbar\omega}-1}.$ (87)
The ancilla is in a generic mixed state,
$\hat{\rho}_{A}(0)=\sum_{n,m}\rho_{A}^{nm}(0){\left|{n}\right\rangle}{\left\langle{m}\right|},$ (88)
so that the total system–environment state at $t=0$ is
$\hat{\rho}(0)=\hat{\rho}_{A}(0)\otimes\bigotimes_{r}\hat{\rho}_{B,r}.$ (89)
To describe a particular stochastic realization, one samples the initial state of each sub-environment from the distribution $P(\alpha_{r,l})$:
$\hat{\rho}(0)=\sum_{n,m}\rho_{A}^{nm}(0)\,{\left|{n}\right\rangle}{\left\langle{m}\right|}\otimes\bigotimes_{r}\bigotimes_{l}{\left|{\alpha_{r,l}}\right\rangle}{\left\langle{\alpha_{r,l}}\right|}.$ (90)
From this point onward, we work in the interaction picture. The corresponding unitary evolution operator factorizes as
$\hat{U}_{I}(t)=e^{-it\sum_{r}\hat{H}_{I,r}}=\sum_{n=0}^{\infty}{\left|{n}\right\rangle}{\left\langle{n}\right|}\otimes\bigotimes_{r}e^{-in\gamma_{r}t\hat{B}_{r}}.$ (91)
Applying this unitary to the initial state gives the evolved density operator
$\hat{\rho}(t)=\sum_{n,m}\rho_{A}^{nm}(0)\,{\left|{n}\right\rangle}{\left\langle{m}\right|}\otimes\bigotimes_{r}\bigotimes_{l}{\left|{\alpha_{r,l}e^{-in\gamma_{r,l}t}}\right\rangle}{\left\langle{\alpha_{r,l}e^{-im\gamma_{r,l}t}}\right|}.$ (92)
In particular, the diagonal sector takes the form of a SBS state:
$\hat{\rho}(t)=\sum_{n}\rho_{A}^{nn}(0)\,{\left|{n}\right\rangle}{\left\langle{n}\right|}\otimes\bigotimes_{r}\underbrace{\bigotimes_{l}{\left|{\alpha_{r,l}e^{-it(\omega_{r,l}+\gamma_{r,l}n)}}\right\rangle}{\left\langle{\alpha_{r,l}e^{-it(\omega_{r,l}+\gamma_{r,l}n)}}\right|}}_{\equiv\,\rho_{r}^{(n)}}+\text{off-diagonal terms},$ (93)
where the diagonal terms of the dynamics forms a spectrum broadcast structure (SBS).
The off-diagonal elements of the ancilla density matrix decay over time due to the interaction with the sub-environments. This decay is quantified by the partial trace over the environment
$\hat{\rho}_{A}=\sum_{n,m}\rho_{A}^{nm}\prod_{r}\prod_{l}\braket{\alpha_{r,l}e^{-it\gamma_{r,l}n}|\alpha_{r,l}e^{-it\gamma_{r,l}m}},$ (94)
the off-diagonal terms $m\neq n$ in the previous expression are products of complex factors of modulus strictly smaller than $1$, whenever $\alpha_{r,l}$ is nonzero. The number of such factors in the product scales with the number of mode in each subenvironment, and the number of environment. Pure dephasing is therefore complete as long as the subenvironments are macroscopic enough. Conversely, in the diagonal terms $n=m$, the products over $r$ and $l$ are equal to unity.
To be able to distinguish between outcomes the trace
$\text{Tr}\{\rho_{r}^{(n)}\rho_{r}^{(m)}\}=\prod_{l}\text{Tr}\{\rho_{r,l}^{(n)}\rho_{r,l}^{(m)}\}$ (95)
should be much smaller than 1. Each factor in the product can be expressed in terms of the coherent-state overlap as
[…]
where $\alpha_{r,l}$ has been considered to be real. Since the total trace is a product over many such factors $l$, the overall overlap $\text{Tr}\{\rho_{r}^{(n)}\rho_{r}^{(m)}\}$ becomes exponentially small as the number of modes increases.